# An Empirical Study Into What Matters for Calibrating Vision–Language Models

### v1

@@ Abstract
Vision–Language Models (VLMs) have emerged as the dominant approach for zero-shot recognition, adept at handling diverse scenarios and significant distribution changes. However, their deployment in risk-sensitive areas requires a deeper understanding of their uncertainty estimation capabilities, a relatively uncharted area. In this study, we explore the calibration properties of VLMs across different architectures, datasets, and training strategies. In particular, we analyze the uncertainty estimation performance of VLMs when calibrated in one domain, label set or hierarchy level, and tested in a different one. Our findings reveal that while VLMs are not inherently calibrated for uncertainty, temperature scaling significantly and consistently improves calibration, even across shifts in distribution and changes in label set. Moreover, VLMs can be calibrated with a very small set of examples. Through detailed experimentation, we highlight the potential applications and importance of our insights, aiming for more reliable and effective use of VLMs in critical, real-world scenarios.
@@ 1 Introduction
In this work, we investigate which factors affect the calibration of VLMs. Starting from prior research (Tu et al., 2023) that demonstrates zero-shot CLIP can be well-calibrated with simple temperature scaling under distribution shifts, we extend this analysis to other CLIP variants and exemplar vision–language models. We then examine whether such a property persists when the calibration dataset varies in (1) distribution, (2) label set (e.g., CIFAR-10 vs. ImageNet), (3) hierarchy level (e.g., “Spider” vs.Black widow”), (4) the number of images, and (5) feature-space distance of calibration set with respect to the target test set.
@@
  • VLMs do not exhibit better uncertainty estimation performance than other models, but become better calibrated than the other classes of models after calibration with temperature scaling – a single parameter variant of Platt Scaling (Platt et al., 1999);
@@
  • VLMs require very few (i.e., less than $100$) samples for calibration;
@@ 2 Related Work
have demonstrated strong capabilities by leveraging web-scale datasets and language supervision to learn joint image–text representations (Bommasani et al., 2021; Radford et al., 2021; Jia et al., 2021). A seminal work by Radford et al. (2021) introduced CLIP, a large VLM trained on $400$ million filtered web-crawled image–text pairs, which exhibits unprecedented zero-shot ability on numerous downstream visual tasks. Inspired by CLIP, various algorithms have been created to enhance the performance of the model (Singh et al., 2022; Li et al., 2022; Zhai et al., 2023; Li et al., 2023a). For example, Li et al. (2022) propose BLIP, a new pre-training framework, that bootstraps captions to effectively leverage noisy web data. Zhai et al. (2023) designs a simple pairwise sigmoid loss which solely operates on the image–text pairs without requiring the global view of pairwise similarities for normalization.
Encouraged by the strong generalizability of VLMs, researchers sought to understand their properties from diverse angles, such as robustness and bias (Fang et al., 2022; Qiu et al., 2022; Tu et al., 2023). For instance, Schiappa et al. (2022) and Qiu et al. (2022) investigate their robustness through perturbations. Fang et al. (2022) point out that the remarkable robustness of CLIP comes from its diverse training distribution. Yuksekgonul et al. (2022) and Thrush et al. (2022) assess the capability of VLMs for encoding compositional information. Liang et al. (2022) study modality gap from views of model initialization and contrastive learning optimization. This paper analyzes the uncertainty estimates of VLMs and considers different aspects that may influence the calibration performance. We also show the usefulness of our observations in a real-world problem setup.
Uncertainty estimation
@@ 3 Definition and Notation
Let $\mathcal{Y}={1,\ldots,K}$ and $\mathcal{X}=\mathbb{R}^{d}$ denote label and input spaces, respectively. A sample $(\mathbf{x},y)$ from an unknown distribution in $\mathcal{X}\times\mathcal{Y}$ is input to a neural network classifier $\mathbf{f}:\mathcal{X}\rightarrow\Delta^{K}$. This classifier outputs a probability distribution over $K$ classes for $\mathbf{x}$, where $\Delta^{K}$ is the $K-1$ dimensional simplex. We assume $\mathbf{f}$ combines two functions: $\mathbf{f}\eqqcolon\mathbf{\sigma}\circ\mathbf{g}$, where $\mathbf{g}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ is a non-probabilistic $n$-way classifier, and $\mathbf{\sigma}:\mathbb{R}^{n}\to\Delta_{n}$ is the softmax operator $\mathbf{\sigma}_{i}(\mathbf{z})=\frac{\exp(\mathbf{z}_{i})}{\sum_{j=1}^{n}\exp(\mathbf{z}_{j})}$ for $i\in\mathcal{Y}$. The output $\mathbf{g}(\mathbf{x})$ is referred to as the logits of $\mathbf{x}$ relative to $\mathbf{f}$. For any input instance $\mathbf{x}$, $\mathbf{f}$ assigns the predicted label $\hat{y}\eqqcolon\arg\max_{i}\mathbf{f}_{i}(\mathbf{x})$ and the corresponding confidence score $\hat{p}\eqqcolon\max_{i}\mathbf{f}{i}(\mathbf{x})$.
@@ 4 Experimental Setup
We consider $35$ vision–language models. These models consist of exemplar VLMs, such as CLIP (Radford et al., 2021), Flava (Singh et al., 2022) and BLIP (Li et al., 2022). In particular, we evaluate zero-shot CLIP models that are trained on different training distributions, such as the WIT (Radford et al., 2021) and LAION (Gadre et al., 2023) datasets, diverse dataset quantities from $3$ million to $2$ billion, and curated training datasets (Xu et al., 2023). CLIP variants with different image encoders are also assessed, including ViT (Dosovitskiy et al., 2020) and ConvNeXt (Liu et al., 2022) encoders, as well as CLIP variants that modify the training objective, such as SigLIP (Zhai et al., 2023), or the training strategy (Li et al., 2023b). Unless specified, for each model, we use their default prompt sets from Radford et al. (2021).
@@
We compare VLM calibration performance with models trained on ImageNet to show that VLMs are well-calibrated despite the distribution shift after temperature scaling. We consider convolutional neural networks, such as ResNet (He et al., 2016) and ConvNeXt (Liu et al., 2022), and vision transformers, exemplified by ViT (Dosovitskiy et al., 2020) and Swin (Liu et al., 2021). These models are trained solely on ImageNet (Deng et al., 2009) or pre-trained on a significantly larger dataset (e.g., ImageNet-21K (Ridnik et al., 2021)). All mentioned models are publicly accessible through OpenCLIP (Ilharco et al., 2021) and TIMM (Wightman, 2019).
@@
We evaluate the calibration of VLM on three standard image classification benchmarks: ImageNet (Deng et al., 2009), CIFAR-10 (Krizhevsky et al., 2009) and DomainNet (Peng et al., 2019). Following the protocol in (Gupta et al., 2021), we divide the validation set of ImageNet into two halves: one for the in-distribution (ID) test set, and the other for learning calibration methods. OOD test sets are ImageNet-V2-A (Recht et al., 2019), ImageNet-R(endition) (Hendrycks et al., 2021), ImageNet-S(ketch) (Wang et al., 2019), and ObjectNet (Barbu et al., 2019). For CIFAR-10, its validation set is used for model calibration, and CIFAR-10.1, CIFAR-10.2 (Recht et al., 2018) and CINIC (Darlow et al., 2018) are used for evaluation. The DomainNet benchmark utilizes the ‘Real’ domain for calibration and evaluates on ‘Painting’ and ‘Sketch’ domains. Note that ImageNet-R and ObjectNet use a reduced subset of classes; we follow the literature (Bello et al., 2021) to select subset of logits for these classes before evaluation.
@@ 5.2 Calibration Across Semantic Hierarchy Levels
Figure 4: Data-efficiency of VLM calibration across diverse datasets. This figure displays the ECE of VLMs as a function of the calibration set size across four datasets: ImageNet-V2-A, ImageNet-S, CINIC, and DomainNet. The ECE values, averaged over ten random seeds, plateau after including merely 4040404050505050 images in the calibration set, closely approximating the error obtained using the full set. This trend is observed despite the high number of classes in DomainNet and ImageNet, where many classes may not be represented even in the calibration set. These results highlight the data-efficiency of VLM calibration.
@@ 5.3 Effect of the Calibration Set Size
Figure 4 plots the calibration error of VLMs with respect to the size of the calibration set for four different target datasets. The calibration error plateaus after $40$-$50$ images and is already close to the error after calibration with the entire set. Notably, DomainNet and ImageNet are $345$-way and $1000$-way classification tasks, and so many classes do not have a single image present in the calibration set. This suggests that calibrating VLMs is very data-efficient: it is a low-dimensional problem.
@@ 5.6 Impact of the Calibration Method
In addition to temperature scaling, we have undertaken an evaluation of another post-hoc calibration method: spline (Gupta et al., 2021). Spline calibration involves deriving a function through spline-fitting, directly aligning classifier outputs with calibrated probabilities. In this section, we investigate whether the observed uncertainty estimation properties of VLMs hold in the spline setting.
In Figure 7, we present our findings, which indicate that our observations regarding the use of temperature scaling for calibration extend to splines as well. For instance, we can observe that when calibrating VLMs using a dataset with different label sets or hierarchy levels than the target test set, similar trends and outcomes persist. This consistency across different post-hoc calibration methods reinforces the robustness of our observations.
@@ 6 Application: Calibration-by-Synthesis
Label set
MECE
MAE
$\rho$
MECE
MAE
$\rho$
Entity13
$0.15$
$15.21$
$0.77$
$0.05$
$5.07$
$0.93$
Entity30
$0.16$
$16.12$
$0.75$
$0.04$
$3.57$
$0.94$
Living17
$0.22$
$21.90$
$0.74$
$0.12$
$11.45$
$0.94$
Nonliving26
$0.13$
$12.73$
$0.79$
$0.05$
$5.29$
$0.91$
Average
$0.17$
$16.49$
$0.76$
$0.07$
$6.35$
$0.93$
Table 1: Uncertainty estimation performance on ImageNet-S for calibration-by-synthesis. We report the mean expected calibration error (MECE), the mean absolute error (MAE) of the average confidence with respect to the model accuracy, and Spearman’s rank correlation ($\rho$). We see that by calibrating on a synthetic set with fine-grained classes, prediction probabilities are informative of the model’s performance and its ranking.

### v2

@@ Abstract
Vision–Language Models (VLMs) have emerged as the dominant approach for zero-shot recognition, adept at handling diverse scenarios and significant distribution changes. However, their deployment in risk-sensitive areas requires a deep understanding of their uncertainty estimation capabilities, a relatively uncharted area. In this study, we explore the calibration properties of VLMs across different architectures, datasets, and training strategies. In particular, we analyze the uncertainty estimation performance of VLMs when calibrated in one domain, label set or hierarchy level, and tested in a different one. Our findings reveal that while VLMs are not inherently calibrated for uncertainty, temperature scaling significantly and consistently improves calibration, even across shifts in distribution and changes in label set. Moreover, VLMs can be calibrated with a very small set of examples. Through detailed experimentation, we highlight the potential applications and importance of our insights, aiming for more reliable and effective use of VLMs in critical, real-world scenarios.
@@ 1 Introduction
We investigate which factors affect the calibration of VLMs. Starting from prior research that demonstrates zero-shot CLIP can be well-calibrated with simple temperature scaling under distribution shifts (Tu et al., 2023), we extend this analysis to other CLIP variants and exemplar vision–language models. We then examine whether such a property persists when the calibration dataset varies in (1) distribution, (2) label set (e.g., CIFAR-10 vs.. ImageNet), (3) hierarchy level (e.g., “Spidervs..Black widow”), (4) the number of images, and (5) feature-space distance of calibration set with respect to the target test set.
@@
  • After calibrating all models with temperature scaling (Platt et al., 1999), vision–language models are better calibrated than the other models in this study, which is not necessarily the case prior to calibration;
@@
  • VLMs require a few samples for calibration. For example, VLMs can be calibrated using temperature scaling, spline fitting (Gupta et al., 2021) or histogram binning (Zadrozny & Elkan, 2001) with less than $100$ samples;
@@ 2 Related Work
have demonstrated strong capabilities by leveraging web-scale datasets and language supervision to learn joint image–text representations (Bommasani et al., 2021; Radford et al., 2021; Jia et al., 2021). A seminal work by Radford et al. (2021) introduced CLIP, a large VLM trained on $400$ million filtered web-crawled image–text pairs, which exhibits unprecedented zero-shot ability on numerous downstream visual tasks. Inspired by CLIP, various algorithms have been created to enhance the performance of the model (Singh et al., 2022; Li et al., 2022; Zhai et al., 2023; Li et al., 2023b). For example, Li et al. (2022) propose BLIP, a new pre-training framework, that bootstraps captions to effectively leverage noisy web data. Zhai et al. (2023) designs a simple pairwise sigmoid loss which solely operates on the image–text pairs without requiring the global view of pairwise similarities for normalization.
Encouraged by the strong generalizability of VLMs, researchers have explored their properties from diverse perspectives, such as robustness and bias. For instance, Schiappa et al. (2022) and Qiu et al. (2022) investigate their robustness through perturbations, while Fang et al. (2022) attribute the remarkable robustness of CLIP to its diverse training distribution. Additionally, Yuksekgonul et al. (2022) and Thrush et al. (2022) assess the capability of VLMs to encode compositional information. Liang et al. (2022) study the modality gap from the perspectives of model initialization and contrastive learning optimization. While Tu et al. (2023) focus on the calibration of CLIP models using temperature scaling on ImageNet, our study extends this to a diverse array of VLMs and various calibration methods (e.g., Spline, histogram binning, and vector scaling). We provide an in-depth analysis of calibration factors, examine the uncertainty estimates of VLMs, and consider different aspects that may influence calibration performance. Furthermore, we demonstrate the practical utility of our findings in a real-world problem setup.
Confidence calibration
@@ 3 Definition and Notation
Let $\mathcal{Y}={1,\ldots,K}$ and $\mathcal{X}=\mathbb{R}^{d}$ denote label and input spaces, respectively. A sample $(\mathbf{x},y)$ from an unknown distribution in $\mathcal{X}\times\mathcal{Y}$ is input to a neural network classifier $\mathbf{f}:\mathcal{X}\rightarrow\Delta^{k}$. This classifier outputs a probability distribution over $k$ classes for $\mathbf{x}$, where $\Delta^{k}$ is the $k-1$ dimensional simplex. We assume $\mathbf{f}$ combines two functions: $\mathbf{f}\eqqcolon\mathbf{\sigma}\circ\mathbf{g}$, where $\mathbf{g}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ is a non-probabilistic $k$-way classifier, and $\mathbf{\sigma}:\mathbb{R}^{k}\to\Delta^{k}$ is the softmax operator $\mathbf{\sigma}_{i}(\mathbf{z})=\frac{\exp(\mathbf{z}_{i})}{\sum_{j=1}^{k}\exp(\mathbf{z}_{j})}$ for $i\in\mathcal{Y}$. The output $\mathbf{g}(\mathbf{x})$ is referred to as the logits of $\mathbf{x}$ relative to $\mathbf{f}$. For any input instance $\mathbf{x}$, $\mathbf{f}$ assigns the predicted label $\hat{y}\eqqcolon\arg\max_{i}\mathbf{f}_{i}(\mathbf{x})$ and the corresponding confidence score $\hat{p}\eqqcolon\max_{i}\mathbf{f}{i}(\mathbf{x})$.
@@ 4 Experimental Setup
We consider $35$ vision–language models. These models consist of exemplar VLMs, such as CLIP (Radford et al., 2021), Flava (Singh et al., 2022) and BLIP (Li et al., 2022). In particular, we evaluate zero-shot CLIP models that are trained on different training distributions, such as the WIT (Radford et al., 2021) and LAION (Gadre et al., 2023) datasets, diverse dataset quantities from $3$ million to $2$ billion, and curated training datasets (Xu et al., 2023). CLIP variants with different image encoders are also assessed, including ViT (Dosovitskiy et al., 2020) and ConvNeXt (Liu et al., 2022) encoders, as well as CLIP variants that modify the training objective, such as SigLIP (Zhai et al., 2023), or the training strategy (Li et al., 2023c). Unless specified, for each model, we use their default prompt sets from Radford et al. (2021).
@@
We compare VLM calibration performance with models trained on ImageNet to show that VLMs are well-calibrated despite the distribution shift after temperature scaling. We consider convolutional neural networks, such as ResNet (He et al., 2016) and ConvNeXt (Liu et al., 2022), and vision transformers, exemplified by ViT (Dosovitskiy et al., 2020) and Swin (Liu et al., 2021). These models are trained solely on ImageNet (Deng et al., 2009) or pre-trained on a significantly larger dataset (e.g., ImageNet-21K (Ridnik et al., 2021)). All the modelsmentioned are publicly accessible through OpenCLIP (Ilharco et al., 2021) and TIMM (Wightman, 2019).
@@
We evaluate the calibration of VLM on three standard image classification benchmarks: ImageNet (Deng et al., 2009), CIFAR-10 (Krizhevsky et al., 2009) and DomainNet (Peng et al., 2019). Following the protocol in (Gupta et al., 2021), we divide the validation set of ImageNet into two halves: one for the in-distribution (ID) test set, and the other for learning calibration methods. OOD test sets are ImageNet-V2-A (Recht et al., 2019), ImageNet-R(endition) (Hendrycks et al., 2021), ImageNet-S(ketch) (Wang et al., 2019), and ObjectNet (Barbu et al., 2019). For CIFAR-10, its validation set is used for model calibration, and CIFAR-10.1, CIFAR-10.2 (Recht et al., 2018b) and CINIC (Darlow et al., 2018) are used for evaluation. The DomainNet benchmark utilizes the ‘Real’ domain for calibration and evaluates on ‘Painting’ and ‘Sketch’ domains. Note that ImageNet-R and ObjectNet use a reduced subset of classes; we follow the literature (Bello et al., 2021) to select subset of logits for these classes before evaluation.
@@ 5.2 Calibration Across Semantic Hierarchy Levels
Figure 4: Data-efficiency of VLM calibration across diverse datasets. This figure displays the ECE of VLMs as a function of the calibration set size across four datasets: ImageNet-V2-A, ImageNet-S, CINIC, and DomainNet. The green stars are the average ECE of calibrated models trained on the dataset. The blue solid and green dashed horizontal lines represent the average ECE before calibration of VLMs and non-VLMs, respectively. The ECE values, averaged over ten random seeds, plateau after including merely 4040404050505050 images in the calibration set, chosen at random, 10 times. The results closely approximate the error obtained using the full set. This trend is observed despite the high number of classes in DomainNet and ImageNet, where many classes may not be represented even in the calibration set. These results highlight the data-efficiency of VLM calibration.
@@ 5.3 Effect of the Calibration Set Size
Figure 4 plots the calibration error of VLMs with respect to the size of the calibration set for four different target datasets. The calibration error plateaus after $40$-$50$ images and is already close to the error after calibration with the entire set. Notably, DomainNet and ImageNet are $345$-way and $1000$-way classification tasks, and so many classes do not have a single image present in the calibration set. This suggests that calibrating VLMs is very data-efficient, and that it is a low-dimensional problem.
@@ 5.5 Transferability of Calibration Across Prompts
Figure 8: Uncertainty estimation performance of VLMs with histogram binning and vector scaling across two factors: calibration set size (up) and prompt sets (bottom). The target domain is ImageNet-S and the calibration set is ImageNet-Validation.
@@ 5.6 Impact of the Calibration Method
In addition to temperature scaling, we undertake an evaluation of three post-hoc calibration methods: spline fitting (Gupta et al., 2021), histogram binning (Zadrozny & Elkan, 2001) and vector scaling (Guo et al., 2017). Spline calibration involves deriving a function through spline fitting (SF), directly aligning classifier outputs with calibrated probabilities. Histogram binning (HB) and vector scaling (VS) both use class-specific parameters and assume an identical label set between calibration set and target set. Hence, HB and VS do not support calibration across label sets and label hierarchy levels. Additionally, as suggested by (Guo et al., 2017), VS requires the presence of samples for each class, meaning that VS demands more images for calibration. This section investigates whether the observed calibration properties of VLMs hold using SF, HB and VS.
In Figure 7 and Figure 8, we present our findings on SF, HB, and VS calibration methods. First, we observe that VLMs achieve lower ECE than calibrated non-VLMs using the three calibration methods under a distribution shift. Second, both SF and HB effectively calibrate VLMs using a small number of samples (i.e. $100$). In contrast, VS requires a larger number of images ($10\%$ of $50,000$ samples) for effective calibration, consistent with the findings in Guo et al. (2017). Third, similar to temperature scaling, both SF and VS can calibrate VLMs using a single prompt, which can differ from the test prompt set. While HB effectively reduces the overall average ECE, it may not benefit all outliers. Lastly, when calibrating VLMs using datasets with different label sets or hierarchy levels compared to the target test set, spline calibration exhibits trends and outcomes similar to those of temperature scaling.
@@ 6 Application: Calibration-by-Synthesis
Label set MECE MAE $\rho$ MECE MAE $\rho$
Entity13 $0.15$ $15.21$ $0.77$ $0.05$ $5.07$ $0.93$
Entity30 $0.16$ $16.12$ $0.75$ $0.04$ $3.57$ $0.94$
Living17 $0.22$ $21.90$ $0.74$ $0.12$ $11.45$ $0.94$
Nonliving26 $0.13$ $12.73$ $0.79$ $0.05$ $5.29$ $0.91$
Average $0.17$ $16.49$ $0.76$ $0.07$ $6.35$ $0.93$
Table 1: Uncertainty estimation performance on ImageNet-S for calibration-by-synthesis. We report the mean expected calibration error (MECE $\downarrow$), the mean absolute error (MAE $\downarrow$) of the average confidence with respect to the model accuracy, and Spearman’s rank correlation ($\rho$uparrow). We see that by calibrating on a synthetic set with fine-grained classes, prediction probabilities are informative of the model’s performance and its ranking.
@@ 7 Conclusion and Discussion
This work leaves open many interesting directions for future research. We primarily focus on calibrating classification confidence estimates for vanilla/backbone VLM networks. It would be interesting to explore the calibration of regression confidence estimates. Some more parameter-efficient fine-tuning methods for prompts can be included to study the transferability of calibration across prompts. Additionally, a broader class of large language model-based VLMs, such as LLaVA (Liu et al., 2023), can be further evaluated and examined. These models could potential present a unique calibration property for the lack of inherent capability to generate classification confidence scores.
Lastly, our scope is the calibration of classification confidence estimates for vanilla/backbone VLM networks, rather than the calibration of regression confidence estimates used for other tasks (e.g., object detection and segmentation). In the context of VLMs, these tasks require compound networks with additional modules. We acknowledge that the findings in this study may not transfer to these other tasks, and that this would be interesting to study in future work.
Acknowledgements
We sincerely thank all the anonymous reviewers and area chairs for their constructive comments and valuable suggestions, which have greatly helped in enhancing this paper.
@@ Acknowledgements
Appendix A Appendix.
A.1 Datasets:
ImageNet (Deng et al., 2009) (https://www.image-net.org/);
ImageNet-V2 (Shankar et al., 2021) (https://github.com/modestyachts/ImageNetV2);
ImageNet-Rendition (Hendrycks et al., 2021) (https://github.com/hendrycks/robustness);
ImageNet-Sketch (Wang et al., 2019) (https://github.com/HaohanWang/ImageNet-Sketch);
ObjectNet (Barbu et al., 2019) (https://objectnet.dev/download.html);
CIFAR-10 (Krizhevsky et al., 2009): (https://www.cs.toronto.edu/ kriz/cifar.html);
CIFAR-10.1 (Recht et al., 2018a): (https://github.com/modestyachts/CIFAR-10.1);
CIFAR-10.2 (Recht et al., 2018a): (https://github.com/modestyachts/CIFAR-10.1);
CINIC (Darlow et al., 2018): (https://www.v7labs.com/open-datasets/cinic-10);
DomainNet (Peng et al., 2019): (http://ai.bu.edu/M3SDA/);
A.2 Models Included in Experiments
(1) Vision–language models:
(1) we use the zero-shot CLIP models provided in OpenCLIP (Ilharco et al., 2021). They are listed as follows in the pattern (architecture, source):
(ViT-B-32openai)
(ViT-B-32laion400m_e32)
(ViT-B-32laion2b_s34b_b79k)
(ViT-B-16openai)
(ViT-B-16laion400m_e32)
(ViT-L-14openai)
(ViT-L-14laion400m_e32)
(ViT-H-14laion2b_s32b_b79k)
(ViT-g-14laion2b_s34b_b88k)
(ViT-bigG-14laion2b_s39b_b160k)
(convnext_baselaion400m_s13b_b51k)
(convnext_base_wlaion_aesthetic_s13b_b82k)
(convnext_xxlargelaion2b_s34b_b82k_augreg)
(ViT-B-32Model-B-32_Data-80M_Samples-34B_lr-1e-3_bs-88k.pt)
(ViT-B-16Model-B-16_Data-80M_Samples-34B_lr-1e-3_bs-88k.pt)
(ViT-L-14Model-L-14_Data-80M_Samples-34B_lr-1e-3_bs-88k.pt)
(ViT-B-32datacomp_m_s128m_b4k)
(ViT-B-32datacomp_s_s13m_b4k)
(ViT-B-16datacomp_l_s1b_b8k)
(ViT-L-14datacomp_xl_s13b_b90k)
(2) Other vision–language models:
(EVA02-B-16merged2b_s8b_b131k)
(EVA02-L-14merged2b_s8b_b131k)
(ViT-B-32-quickgelumetaclip_400m)
(ViT-B-16-quickgelumetaclip_fullcc)
(ViT-B-16-SigLIPwebli)
(ViT-B-16-SigLIP-256webli)
(ViT-H-14-CLIPAdatacomp1b)
(ViT-L-14-CLIPAdatacomp1b)
Flava can be derived from Hugging Face, Transformer module (Wolf et al., 2020).
BLIP (Li et al., 2022), BLIP-2 (Li et al., 2023b) and ALBEF (Li et al., 2021) models are publicly available through LAVIS package (Li et al., 2023a).
(2) ImageNet-trained models:
we compare VLMs’ uncertainty estimate performance with following models from TIMM (Wightman, 2019)
Trained from scratch on ImageNet:
swin_base_patch4_window12_384
deit_small_distilled_patch16_224
deit_base_patch16_224
swin_small_patch4_window7_224
deit_base_patch16_384
wide_resnet50_2
convnext_base
tv_resnet50
densenet121
inception_v4
resmlp_36_224
xception
vgg19
Pre-trained on a larger dataset then fine-tuned on ImageNet training split:
vit_base_patch16_224.orig_in21k_ft_in1k
vit_base_patch32_224
vit_large_r50_s32_384
beit_large_patch16_224
vit_large_patch16_384
swin_large_patch4_window12_384
convnext_small.in12k_ft_in1k
ig_resnext101_32x16d
mixer_b16_224_miil
resnetv2_50x1_bitm
ig_resnext101_32x8d
A.3 Prompt Template
In this section, we provide the prompt set used in our experiments. For experiments included in Section 5.1, Section 5.2, Section 5.4 and Section 5.6, we also use default prompt sets provided by Radford et al. (2021). For experiments in Section 5.5, the calibration prompt set is “a photo of a $<$class$>$”. In Section 6, the calibration prompt set is “a photo of a $<$class$>$”, while the test prompt set is generated by CuPL.
A.4 Computation Resources
PyTorch version is 1.10.0+cu111 and timm version is 0.8.21dev0. All experiment is run on one 3090 and the CPU AMD EPYC 7343 16-Core Processor.
A.5 Experiments on Specialized Domain: Satellite Images
We further conduct experiments on whether our findings maintain on more specialized domains: Satellite images. We use EuroSAT (Helber et al., 2019) as the target set. Two different calibration datasets are considered: 1) ImageNet validation set; 2) A generated dataset which consists of $68$ images from finer-grained classes of Living17 hierarchy. The calibration method is temperature scaling. We see that our observations persist on such a specialized domain that VLMs can be calibrated across different label sets using a few samples.
Figure 10: Uncertainty estimate performance of VLMs when the target domain contains Satellite images. Left: The calibration set is ImageNet-Validation set. Right: The calibration set is a synthesized dataset containing fine-grained classes of Living17.
A.6 Additional Results for Adaptability to Different Calibration Label Sets
Following the same practice as Section 5.1, we additionally validate that the findings on ImageNet-S and ImageNet-R persist when the target datasets change to ImageNet-A or ObjectNet.
Figure 11: Adaptability of VLMs to different calibration label sets. Left: Calibration error reduction. Here, we observe a significant decrease in the expected calibration error for VLMs following cross-label-set calibration, as opposed to when no calibration is applied. Right: Correlation between VLM prediction probability and classification accuracy. This graph illustrates the classification accuracy of VLMs on ObejctNet and ImageNet-A against their average prediction probability, before and after calibration with CIFAR-10-Val or DomainNet-Real. Each point represents a model, with the dashed black line indicating perfect calibration ($y=x$). The data showcases a strong linear and rank correlation, even when models are calibrated on label sets different from the target, proving the effectiveness of cross-label-set calibration for VLMs.
A.7 Additional Results for Calibration Across Semantic Hierarchy Levels
We also show the ECE of VLMs calibrated with label hierarchies differing in granularity from the target dataset (ImageNet-V2-A). The results validate that VLMs can be calibrated across label hierarchies.
Figure 12: Robustness of VLM calibration to label hierarchy levels. This figure presents box plots summarizing the calibration errors (ECEs) of VLMs calibrated with label hierarchies differing in granularity from the target dataset (ImageNet-V2-A). The top row shows calibration at a coarser level, and the bottom row at a finer level. Despite not matching the calibration precision of same-level calibration, the minimal differences indicate the robustness of VLM calibration to label granularity.
A.8 Transferability of Calibration Across Prompts Using Advanced Prompt Tuning Methods
We additionally test whether our observations that VLMs can be calibrated by using “a photo of a $<$class$>$” persist when the test prompt set is obtained via advanced prompt fine-tuning methods, such as MaPLe (Khattak et al., 2023a) or PromptSRC (Khattak et al., 2023b). We use ViT-B/16 as the visual backbone and fine-tune them on ImageNet validation with $16$ shots per class. Following the procedure in 5.5, we randomly selected a single prompt for calibration and evaluated whether the searched temperature scalar is useful on ImageNet-Sketch, ImageNet-V2 and ImageNet-Rendition for the prompt-tuned models. The results are summarized in . We have a consistent observation with 6 that VLM can be calibrated with one single prompt.
ImageNet-S ImageNet-V2-A ImageNet-R
Before After Before After Before After
MaPLe $0.0364$ $0.0073$ $0.0159$ $0.0165$ $0.0264$ $0.0073$
PromptSRC $0.0323$ $0.0071$ $0.0175$ $0.0173$ $0.0328$ $0.0071$
Table 2: Uncertainty estimation performance on ImageNet-S, ImageNet-V2-A and ImageNet-R for cross prompt set calibration. We report the expected calibration error (ECE $\downarrow$) before and after calibration on ImageNet validation set. We see that by calibrating on one single prompt on ImageNet-Val with temperature scaling, VLMs become well-calibrated.